# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@
Spatial Scale of Agglomeration and Dispersion
@@
Takashi Akamatsu Tomoya Mori Tohoku Univ. Kyoto Univ., RIETI Minoru Osawa Yuki Takayama Kyoto Univ. Science Tokyo TA: akamatsu@plan.civil.tohoku.ac.jp, TM: mori@kier.kyoto-u.ac.jp, MO: osawa.minoru.4z@kyoto-u.ac.jp, YT: takayama.y.cc65@m.isct.ac.jp. We are indebted to Treb Allen, Kristian Behrens, Donald Davis, Gilles Duranton, Yasusada Murata, Jonathan Newton, Michael Pflüger, Diego Puga, Esteban Rossi-Hansberg, Shunsuke Segi, Akihisa Shibata, and Jacques-François Thisse for their constructive comments. We are grateful to anonymous referees and to all seminar and conference participants. This research has been supported by JSPS KAKENHI 17H00987, 18H01556, 19K15108, 21K04299, 22K04353, 23K22880, 23K22887, 24K00999, and 25H00543. This research was conducted as part of the project, “Agglomeration-based Framework for Empirical and Policy Analyses of Regional Economies,” undertaken at the Research Institute of Economy, Trade and Industry. We acknowledge financial support from the Kajima Foundations and the Murata Science Foundation.
(July 4, 2025)
Cities around the world have become fewer, larger, and more spatially dispersed. We offer a theoretical framework that can account for this dual transformation—economy-wide concentration and local spreading—by distinguishing between two types of dispersion forces: “local” forces acting within cities, and “global” forces acting across them. The distinction leads to a systematic classification of spatial models into a few fundamental types, each entailing distinct spatial patterns and comparative statics. The framework serves to reconcile divergent empirical findings and clarifies how spatial outcomes of transport policies may depend on the spatial scale at which relevant dispersion forces operate.
@@
Keywords: agglomeration patterns; spatial scale of dispersion forces, many regions, economic geography
@@ 1   Introduction
Urban development often exhibits patterns that appear contradictory at first glance. In Japan, the post-1960 construction of nationwide high-speed railway and highway networks [^1] Between 1970 and 2020, the total length increased from 1,119 km to 9,050 km for highways, and from 515 km to 3,106 km for high-speed railways—an increase of over eightfold and sixfold, respectively. See Fig. B.1 in Appendix B. took place alongside rapid demographic shifts. Over the subsequent five decades, urban populations became increasingly concentrated in larger cities: the population share of Japan’s top 100 cities rose by 19 %, while that of the remaining cities fell by 17 %, as reflected in the steeper slope of the rank–size plot (Fig. 1A). At the same time, however, population became more spatially dispersed within each city: between 1970 and 2020, the average city experienced a 35 % decline in its maximum population density and a 24 % decline in average density (Fig. 1B). Local spreading is also evident in within-city population distributions (Fig. 1C). The co-occurrence of economy-wide concentration and local spreading has also been observed in countries such as China, France, and the United States over the recent decades. [^2] See Appendix B. Combes-et-al-DP2023 obtained similar results for France for a much longer duration of time by constructing original panel data for French cities during the 1760–2020 period. The recurring dual pattern across varied institutional and historical settings points to a deeper mechanism operating across spatial scales.
@@ A Population rank–size plot
B Maximum and average population density
@@ C The population distribution within Tokyo in 1970 (left) and 2020 (right)
Note: A city is defined as a cluster of contiguous 1km-by-1km grid cells, each with a population density of at least 1,000/km2 and collectively comprising at least 10,000 residents (see Appendix B for details). The number of identified cities decreased from 504 in 1970 to 431 in 2020. Panel (B) shows the annual cross-city arithmetic means of maximum and average population densities, along with the 90th percentile band. Panel (C) visualizes the within-city population distribution in Tokyo; similar figures for Osaka (Japan’s second-largest city) and Nagoya (third-largest) are provided in Figs. 4A and 4B.
@@ F Polycentric distribution
Motivated by multi-scale agglomeration and dispersion phenomena, this study develops a conceptual framework to distinguish between “local” and “global” dispersion forces, each acting at a distinct spatial scale. While many spatial models implicitly embed local or global dispersion forces, we formalize the distinction that has received limited systematic attention in the literature. Our framework follows the literature on spatial agglomeration to employ stylized geographical settings such as a symmetric two-region economy. This allows us to provide systematic comparison of endogenous mechanisms across a wide range of spatial models, including canonical quantitative models such as Redding-Sturm-AER2008 and Allen-Arkolakis-QJE2014.
@@
To develop intuition for the proposed distinction between local and global dispersion forces, consider a hypothetical many-location economy with agents—such as consumers or firms—who make location choices. Suppose there are some agglomeration forces so that agents benefit from proximity to others. In the absence of any countervailing dispersion force, all agents concentrate in a single location (Fig. 1D).
@@
Now, consider “local” dispersion forces: negative externalities that arise within a location, depend on its own population, and affect only its residents. A representative example is crowding in non-tradable sectors, such as housing scarcity. Specifically, if housing supply is inelastic, population growth within a city drives up prices, creating incentives for residents to relocate to nearby, less expensive areas. Agents relocate incrementally toward the periphery to mitigate congestion while still benefiting from proximity to others. In equilibrium, this tension gives rise to a single-peaked, monocentric spatial pattern, with a dominant center surrounded by lower-density fringes, as shown in Fig. 1E.
By contrast, “global” dispersion forces are defined as negative spillovers across locations. A canonical example is the market crowding effect resulting from interregional trade, combined with immobile factors such as land or other natural resources. In this case, interregional trade allows the central location with large agglomeration benefits to outcompete neighboring areas, leading to revenue losses in the local markets of those regions. If transport costs are sufficiently high, producers have an incentive to relocate to more distant areas to escape intense price competition at the center and to better capture local demand, while benefiting from local immobile resources such as land or natural endowments. This creates incentives for firms and workers to spread out across space, giving rise to multiple agglomerations that compete for market access and resources, and thereby spatially repel each other (Fig. 1F).
@@
As transport costs decline, local and global dispersion forces exert countervailing influences. At the global scale, improved transport access extends the spatial reach of firms and consumers, intensifying market crowding. This undermines smaller agglomerations, driving economic activities to concentrate in a limited number of major hubs. At the local scale, however, lower transport costs reduce the relative advantages of proximity, fostering spatial spread within cities due to congestion forces. These opposing mechanisms can jointly give rise to a dual pattern: economy-wide concentration alongside intraurban decentralization.
@@
The concepts of local and global dispersion forces have important implications for spatial policy analysis. This study shows that, while there are various possible microfoundations for these forces, distinguishing between them helps clarify the equilibrium structure and comparative statics of models of economic agglomeration. In particular, we provide a systematic taxonomy of spatial models into three broad model classes: (i) those with only local dispersion forces, (ii) those with only global dispersion forces, and (iii) those that incorporate both. These model classes can imply fundamentally different predictions regarding the spatial impact of declining transport costs. Thus, quantitative models can yield sharply divergent counterfactual predictions depending on the dispersion forces they embed. For example, declining transport costs tend to promote spatial spreading under local dispersion forces (e.g., Helpman-Book1998; Allen-Arkolakis-QJE2014), but foster further agglomeration under global dispersion forces (e.g., Krugman-JPE1991). Accordingly, transport policies intended to support peripheral regions may succeed or backfire, depending on whether the model used to design them adequately represents the dispersion forces relevant to the spatial problem at hand.
@@ 2   The two-region economy
We begin with a parsimonious two-region framework to introduce key concepts, and later discuss richer settings including general equilibrium models (e.g., Krugman-JPE1991; Helpman-Book1998; Tabuchi-JUE1998; Redding-Sturm-AER2008; Allen-Arkolakis-QJE2014). Then, Section 3 will consider a stylized many-region setting.
Consider a closed two-region economy, and let $\mathcal{I}\equiv\{1,2\}$ be the set of regions. There are a continuum of perfectly mobile agents, such as households, who choose their location to maximize utility. Let $x_{i}\geq 0$ denote the mass of agents in region $i$. The spatial distribution is $\bm{x}=(x_{1},x_{2})$. We let $x_{1}+x_{2}=1$ for normalization. Let $v_{i}(\bm{x})$ denote the indirect utility of agents in region $i$ as a function of $\bm{x}$. A spatial equilibrium is a distribution $\bm{x}$ such that no agent has an incentive to relocate. [^3] Formally, $\bm{x}$ is an equilibrium if $\bar{v}\geq v_{i}(\bm{x})$ for all $i$ with equality if $x_{i}>0$, where $\bar{v}$ is an endogenous scalar representing the corresponding equilibrium utility level.
Throughout this section, we consider a series of specifications for the indirect utility function $v_{i}(\bm{x})$ to highlight the difference between “local” and “global” dispersion forces. In doing so, we assume that regional characteristics are symmetric to isolate the trade-offs between agglomeration and dispersion forces driven by transport costs and the endogenous spatial distribution of mobile agents.
Transport across the regions is costly and captured by a proximity matrix $[\phi_{ij}]$, where $\phi_{ij}\in(0,1]$ measures the ease of transport from region $i$ to $j$. Higher values indicate better access. Throughout this section, we assume that
$\begin{bmatrix}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{11}\end{bmatrix}=\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix},$ (1)
where $\phi\in(0,1)$ is the transport cost parameter.
@@ 2.1   The Beckmann model: A “local” dispersion force
We begin with a minimal model that exemplifies a “local” dispersion mechanism. This also serves to revisit the classic intuition of agglomeration as symmetry-breaking.
Consider the following reduced-form model featuring the trade-off between intraregional congestion costs and agents’ desire to be close to each other:
$v_{i}(\bm{x})=x_{i}^{-\beta}\Big{(}\sum_{j\in\mathcal{I}}\phi_{ij}x_{j}\Big{)}^{\alpha}$ $(\alpha,\beta>0).$ (2)
@@
The first term ($x_{i}^{-\beta}$) captures congestion force, while the second reflects proximity-based positive externalities. We refer to this model as the Beckmann model after the seminal work by Beckmann-Book1976.444Beckmann studied spatial agglomeration in a continuous one-dimensional space. We consider its discrete-space and multiplicative version to streamline the exposition.
The uniform distribution of agents $\bar{\bm{x}}=(\tfrac{1}{2},\tfrac{1}{2})$ is an equilibrium, as both regions yield equal utility at $\bar{\bm{x}}$. A crucial feature of the model is that $\bar{\bm{x}}$ may not be stable. If the agglomeration force dominates the dispersion force, a tiny migration shock encourages further migration of agents, pushing the system away from $\bar{\bm{x}}$. Accordingly, a spatial pattern with $x_{1}>x_{2}$ or $x_{1}<x_{2}$ becomes a stable equilibrium. That is, agglomeration of mobile agents occurs without ex-ante regional asymmetries.
To determine the stability of $\bar{\bm{x}}$, we must check the incentive of a hypothetical migrant. Specifically, the utility gain from migration of a marginal mover from region $2$ to $1$ can be measured by the following elasticity of the utility difference:
@@
$\omega$ $\equiv\frac{\bar{x}}{\bar{v}}\times\left.\frac{\partial{\left(v_{1}(\bm{x})-v_{2}(\bm{x})\right)}}{\partial{x_{1}}}\right|_{\bm{x}=\bar{\bm{x}}}=\frac{\bar{x}}{\bar{v}}\times\left.\left(\frac{\partial{v_{1}(\bm{x})}}{\partial{x_{1}}}-\frac{\partial{v_{2}(\bm{x})}}{\partial{x_{1}}}\right)\right|_{\bm{x}=\bar{\bm{x}}},$ (3)
where $\bar{x}\equiv\tfrac{1}{2}$ and $\bar{v}=v_{1}(\bar{\bm{x}})=v_{2}(\bar{\bm{x}})$ is the equilibrium utility level at $\bar{\bm{x}}$. The term $\frac{\bar{x}}{\bar{v}}$ simplifies the resulting expression.
If $\omega<0$, it indicates that a marginal increase in the mass of agents in region $1$ induces a relative utility decrease therein compared to region $2$. Migrants are strictly worse off, implying reverse migration. Hence, $\bar{\bm{x}}$ is stable. If $\omega>0$ instead, migrants are strictly better off. Further migration should occur. That is, $\bar{\bm{x}}$ is unstable.
Evaluating Eq. 3 for the Beckmann model yields
@@
$\omega$ $=-\beta+\alpha\Theta,\quad\text{where}\quad\Theta\equiv\frac{1-\phi}{1+\phi}.$ (4)
The term $\Theta\in(0,1)$ is a convenient index of transport cost. It decreases as $\phi$ increases, with $\Theta\to 1$ as $\phi\to 0$ (regional autarky) and $\Theta\to 0$ as $\phi\to 1$ (frictionless transport).
@@
The expression of $\omega$ is useful in two respects. First, it succinctly captures the trade-off between two opposing forces. The negative term $-\beta$ reflects the dispersion force, whereas the positive term $\alpha\Theta$ reflects the agglomeration force. The stability of $\bar{\bm{x}}$ (i.e., the sign of $\omega$) depends on which force dominates. Specifically,
$\omega\lesseqgtr 0\quad\Leftrightarrow\quad\alpha\Theta\lesseqgtr\beta$ (5)
with the same signs for the inequalities. For example, if congestion force is too strong ($\beta\geq\alpha$), $\omega<0$ for all $\phi\in(0,1)$, implying no agglomeration can occur.
Second, the expression also elucidates the spatial nature of the economic forces. The agglomeration force depends on transport costs through $\Theta$, whereas the dispersion force does not. The dispersion force is strictly “local” in that it is determined solely by the population within each region, is borne only by local residents, and remains unaffected by interregional transport conditions. By contrast, the agglomeration force is “global,” as it incorporates interregional interactions. As distance becomes less costly (as $\Theta$ decreases), the agglomeration force $\alpha\Theta$ weakens, since the benefit of being close to others diminishes. [^5] Note that, in the opposite extreme of $\phi\to 0$, $\Theta\to 1$ and $\alpha\Theta\to\alpha$. Since the regions are in autarky, the agglomeration force reduces to local spillovers à la Henderson-AER1974, and Eq. 5 reduces to whether positive spillovers $\alpha$ is smaller than negative ones $\beta$ or not. If $\beta<\alpha$, transport costs can matter: $\bar{\bm{x}}$ is stable for low transport costs (small $\Theta$Leftrightarrow large $\phi$) but unstable for high transport costs (large $\Theta$Leftrightarrow small $\phi$).

### v8

# Spatial Scale of Agglomeration and Dispersion: Number, Spacing, and the Spatial Extent of Cities

@@
Spatial Scale of Agglomeration and Dispersion:
@@
Takashi Akamatsu Tomoya Mori Tohoku Univ. Kyoto Univ., RIETI Minoru Osawa Yuki Takayama Kyoto Univ. Science Tokyo TA: akamatsu@plan.civil.tohoku.ac.jp, TM: mori@kier.kyoto-u.ac.jp, MO: osawa.minoru.4z@kyoto-u.ac.jp, YT: takayama.y.cc65@m.isct.ac.jp. We are deeply grateful to Co-Editor Gilles Duranton for his guidance and support throughout the review process, which enormously improved the paper. We also thank two thoughtful referees for very helpful comments. We thank Treb Allen, Kristian Behrens, Donald Davis, Yasusada Murata, Jonathan Newton, Michael Pflüger, Diego Puga, Esteban Rossi-Hansberg, Shunsuke Segi, Akihisa Shibata, and Jacques-François Thisse for their comments, including those for the earlier versions of this paper. We also thank all seminar and conference participants. This research has been supported by JSPS KAKENHI 17H00987, 18H01556, 19K15108, 21K04299, 22K04353, 23K22880, 23K22887, 24K00999, and 25H00543. This research was conducted as part of the project, “Agglomeration-based Framework for Empirical and Policy Analyses of Regional Economies,” undertaken at the Research Institute of Economy, Trade and Industry. We acknowledge financial support from the Kajima Foundations and the Murata Science Foundation.
How does transport cost affect the spatial organization of economic activities? This study develops a theoretical framework that distinguishes between two types of dispersion forces in spatial models: “local” dispersion forces acting within cities, and “global” dispersion forces acting across them. The distinction leads to a systematic classification of spatial models into a few fundamental types, each with distinct endogenous spatial patterns and comparative statics in response to changes in transport costs. The framework reconciles empirical findings and clarifies how transport-induced reorganization of economic activities can depend on the spatial scale of dominant dispersion forces.
@@
Keywords: agglomeration; spatial scale; quantitative spatial model.
@@ 1  Introduction
Urban development often exhibits patterns that appear contradictory at first glance. In Japan, over the five decades from 1970 to 2020, urban populations became increasingly concentrated in larger cities. The population share of the top 100 cities rose by 19 %, while that of the remaining cities fell by 17 %, as reflected in the flatter slope of the rank–size plot (Fig.˜1A). Concurrently, however, populations became more spatially dispersed within individual cities. The average city experienced a 35 % decline in its maximum population density and a 24 % decline in average density (Fig.˜1B), a trend evident in within-city distributions (e.g., Fig.˜1C). Such a dual trend of economy-wide concentration and intra-urban decentralization is not unique to Japan but is observed across diverse contexts, including China, France, and the United States. [^1] See Appendix B for more detailed discussion. Combes-et-al-DP2023 report similar evidence for France using newly constructed panel data spanning 1760–2020.
@@ A Population rank–size plot
B Population density (1,000/km2)
@@ C The population distribution within Tokyo in 1970 (left) and 2020 (right)
Note: A city is defined as a cluster of contiguous 1km-by-1km grid cells, each with a population density of at least 1,000/km2 and collectively comprising at least 10,000 residents. The number of cities decreased from 504 in 1970 to 431 in 2020. Panel (B) shows the annual cross-city arithmetic means of maximum and average population densities along with the 95% bootstrap confidence interval. Panel (C) shows the within-city population distribution in Tokyo in 1970 and 2020. For further discussion, see Appendix B.
@@ F Polycentric distribution
How can population simultaneously concentrate across cities and spread out within them? Transport costs have been central to theories of economic agglomeration and dispersion. Indeed, Japan’s spatial reorganization coincided with the rapid expansion of nationwide high-speed railway and highway networks, which were essentially developed from scratch between 1970 and 2020.222Between 1970 and 2020, the total length increased from 1,119  km to 9,050  km for highways, and from 515  km to 3,106  km for high-speed railways. This is an increase of more than eight times and six times, respectively. See Fig. B.1 in Appendix B. However, interpreting this dual pattern through existing theory is not straightforward. This is because most analyses rely on highly stylized few-location settings, , which makes it difficult to analyze how agglomeration and dispersion occur in complex spatial economies with many locations.
To address this, the present study develops a unified theoretical framework that distinguishes between “local” dispersion forces acting within cities and “global” dispersion forces acting across them. The core intuition is that dispersion, or the tendency for economic activity to spread out, operates at two distinct spatial scales: one that pushes agents toward the fringes of their own city, and another that repels cities away from one another. This distinction provides a coherent basis for classifying spatial models, clarifying how their comparative statics differ and how transport-induced reorganization depends on the dominant type of dispersion force in many-location settings.
@@
To understand the core intuition, consider a hypothetical many-location economy with mobile agents, such as households or firms, making location choices. For exposition purposes, “locations” refer to generic discrete units such as regions, counties, cities, or grid cells. Suppose that agents benefit from proximity to others through some agglomeration forces. In the absence of dispersion forces, everyone concentrates in a single location, as illustrated by Fig.˜1D.
@@
First, consider the negative externalities that arise within a location, depend on its own population, and affect only its residents. We term such externalities “local” dispersion forces. A representative example is crowding in the market for non-tradable goods. For instance, if the housing supply is inelastic, population growth within a location drives up housing prices, creating incentives for residents to relocate to nearby, less expensive areas. Agents move incrementally toward the periphery to mitigate congestion while still benefiting from proximity to the core. In equilibrium, this tension gives rise to a single-peaked monocentric spatial pattern, characterized by a dominant center surrounded by lower-density fringes, as Fig.˜1E illustrates.
By contrast, some negative externalities spill across locations. We refer to these as “global” dispersion forces. A key example is market crowding through interregional trade in the presence of immobile factors such as land or other natural resources. Through trade, a large central city can use its agglomeration advantages to dominate nearby markets, making surrounding locations unattractive for producers. When transport costs are sufficiently high, producers may instead locate farther from the center, where competition is weaker and local demand can support entry. The result is a spatial pattern with multiple agglomerations that compete for market access and scarce resources and therefore repel each other across space, as Fig.˜1F illustrates.
@@
As transport costs decline, local and global dispersion forces exert countervailing influences. On a global scale, winners and losers can emerge. Improved transport access extends the spatial reach of firms and consumers, intensifying market crowding between economic centers. This undermines smaller agglomerations, driving economic activities to concentrate further in a limited number of major hubs. However, on the local scale, lower transport costs reduce the relative advantages of proximity, fostering spatial spread within cities due to congestion forces. These opposing mechanisms can jointly give rise to a dual pattern: economy-wide concentration alongside decentralization within each agglomeration.
@@
Distinguishing between local and global dispersion forces helps clarify the equilibrium spatial patterns implied by models of economic agglomeration and the comparative statics they generate. A wide range of spatial models can be classified into three broad types: (i) models with only local dispersion forces, (ii) models with only global dispersion forces, and (iii) models that incorporate both. These model classes yield fundamentally different predictions about how declining transport costs shape the spatial economy. In turn, quantitative spatial models can produce sharply divergent counterfactual outcomes depending on the dispersion forces they embed. For example, lower transport costs tend to promote spatial spread across locations when local dispersion forces dominate (e.g., Helpman-Book1998; Allen-Arkolakis-QJE2014), but promote further agglomeration toward central locations when global dispersion forces are primary (e.g., Krugman-JPE1991). In short, whether transport improvements lead to spatial spreading or further concentration depends on the spatial scale at which dispersion forces operate. Consequently, transport policies intended to support peripheral locations may succeed or backfire depending on whether the models guiding these policies adequately capture the relevant dispersion forces.
@@ 2  The two-region economy
To introduce key concepts, we consider two-location models throughout this section, and proceed to many-location settings in later sections.
Preliminaries. There are perfectly mobile agents (e.g., households) who choose their location to maximize utility. Throughout, locations are called regions for convenience. Let $x_{i}\geq 0$ denote the continuous mass of agents in the region $i\in\{1,2\}$, where $x_{1}+x_{2}=1$. The indirect utility of agents in the region $i$ is a differentiable function of the spatial distribution $\bm{x}=(x_{1},x_{2})$ and is indicated by $v_{i}(\bm{x})$. A spatial equilibrium is a spatial distribution $\bm{x}^{*}$ in which no agent is motivated to relocate. For example, if $x_{1}^{*},x_{2}^{*}>0$, then $\bm{x}^{*}$ is a spatial equilibrium if and only if $v_{1}(\bm{x}^{*})=v_{2}(\bm{x}^{*})$.
Transport between regions is costly, and $\phi\in(0,1)$ measures the ease of transport between regions. Higher values of $\phi$ indicate better access. For later use, we also introduce the proximity matrix $\phi_{ij}]$ following Matsuyama-RIE2017, where $\phi_{ij}\in(0,1]$ is the ease of transport from region $i$ to $j$. For the two-region case in this section,
$\begin{bmatrix}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{22}\end{bmatrix}=\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix}$. (1)
To focus on forces driven by transport costs and the endogenous spatial distribution of agents, region-specific characteristics (e.g., innate amenity or productivity) are assumed to be homogeneous. With these common settings in place, we consider a series of specifications for $\bm{v}(\bm{x})=(v_{1}(\bm{x}),v_{2}(\bm{x}))$, including general equilibrium models such as Krugman-JPE1991; Helpman-Book1998; Tabuchi-JUE1998; Redding-Sturm-AER2008; Allen-Arkolakis-QJE2014, to illustrate the core ideas.
@@ 2.1  The Beckmann model: A “local” dispersion force
We start with the following parsimonious specification:
$v_{i}(\bm{x})=x_{i}^{-\beta}\Big(\sum_{j}\phi_{ij}x_{j}\Big)^{\alpha}$ $(\alpha,\beta>0).$ (2)
@@
The first term, $x_{i}^{-\beta}$, captures the localized congestion force within each region, while the second term is positive externalities between regions or the agents’ desire to be close to each other. We call this the Beckmann model, following Beckmann-Book1976.333Beckmann studied spatial agglomeration in a continuous one-dimensional space. To streamline the exposition, we consider its discrete-space and multiplicative analog.
The symmetric distribution of agents, $\bar{\bm{x}}=(\tfrac{1}{2},\tfrac{1}{2})$, is always a spatial equilibrium. In fact, if we define the utility difference between the two regions by
$\Delta(\bm{x})\equiv v_{1}(\bm{x})-v_{2}(\bm{x}),$ (3)
we confirm $\Delta(\bar{\bm{x}})=0$.
While $\bar{\bm{x}}$ is an equilibrium, with both agglomeration and dispersion forces, $\bar{\bm{x}}$ is not stable if the former dominates the latter. To assess the stability of $\bar{\bm{x}}$, we examine the incentive of movers. Consider a small mass of agents moving from region 2 to 1, so that $x_{1}$ rises and $x_{2}$ falls symmetrically. If $\Delta$ decreases under this perturbation, then $\Delta<0$ after the move: utility in region 2 exceeds that in region 1, and the movers will prefer to return. This induces a restoring force and hence $\bar{\bm{x}}$ is locally stable. If $\Delta$ increases, then $\Delta>0$ after the move: the movers prefer to stay in region 1, and the deviation attracts additional movers now that the utility is higher in region $1$. Hence, $\bar{\bm{x}}$ is unstable.
Figure 3: Stability of the symmetry in the Beckmann model ($\alpha=1/2,\beta=1/6$)
Figure˜3 illustrates the above discussion. The symmetric equilibrium $\bar{\bm{x}}$ is stable in the left panel. In the right panel, $\bar{\bm{x}}$ is unstable and two additional stable spatial equilibria arise, each exhibiting endogenous agglomeration, in which one region becomes larger despite the perfect symmetry of the regional fundamentals.
We can formalize the above discussion by a simple stability criterion. [^4] All omitted proofs are in Appendix A.
Lemma 1.
Given a differentiable $\bm{v}$, define the utility gain for marginal movers by
@@
$\omega$ $\equiv\frac{\bar{x}}{\bar{v}}\cdot\left.\frac{\partial{\Delta(\bm{x})}}{\partial{x_{1}}}\right|_{\bm{x}=\bar{\bm{x}}}=\frac{\bar{x}}{\bar{v}}\cdot\left.\left(\frac{\partial{v_{1}(\bm{x})}}{\partial{x_{1}}}-\frac{\partial{v_{2}(\bm{x})}}{\partial{x_{1}}}\right)\right|_{\bm{x}=\bar{\bm{x}}},$ (4)
where $\bar{x}\equiv\tfrac{1}{2}$ and $\bar{v}=v_{1}(\bar{\bm{x}})=v_{2}(\bar{\bm{x}})>0$. Then, the symmetric equilibrium $\bar{\bm{x}}$ is stable if $\omega<0$, and unstable if $\omega>0$. ∎
In Eq.˜4, normalization by $\bar{x}/\bar{v}$ only simplifies the final expression of $\omega$.
In the Beckmann model, we compute
@@
$\omega$ $=-\beta+\alpha\Theta,\qquad\text{where}\qquad\Theta\equiv\frac{1-\phi}{1+\phi}\in(0,1).$ (5)
@@
The negative term $\beta$ represents the dispersion force, while the positive term $\alpha\Theta$ represents the agglomeration force. The sign of $\omega$, and hence the stability of $\bar{\bm{x}}$, depends on which force dominates: $\bar{\bm{x}}$ is stable if $\beta>\alpha\Theta$, and unstable if $\beta<\alpha\Theta$.
In Eq.˜5, the dispersion force does not depend on $\phi$ because it is not affected by interregional transport conditions. By contrast, the agglomeration force depends on $\phi$ as it incorporates interregional interactions. In particular, $\alpha\Theta$ decreases as $\phi$ increases. That is, the benefit of becoming close to others is small if transport is less costly. [^5] If instead the regions are in autarky ($\phi=0$), the agglomeration force reduces to local spillovers à la Henderson-AER1974. Since $\Theta=1$ and $\alpha\Theta=\alpha$, the stability condition is whether $\alpha<\beta$ or not.
Importantly, $\Theta$ itself is the proximity gain for marginal movers at each $\phi$. To see this, set $\alpha=1$ and $\beta=0$. Then, $v_{i}(\bm{x})=\sum_{j}\phi_{ij}x_{j}$ corresponds to a parsimonious measure of proximity. Since $\omega=\Theta$ in this case, $\Theta$ indeed represents the proximity gain.
With Eq.˜5, we can determine the stability of $\bar{\bm{x}}$ in the Beckmann model. If $\beta\geq\alpha$, congestion force is so strong that $\omega<0$ for all $\phi\in(0,1)$. Agglomeration cannot occur, as $\bar{\bm{x}}$ is always stable. If $0<\beta<\alpha$, the level of $\phi$ matters. As Fig.˜3 illustrates, $\bar{\bm{x}}$ is stable if $\phi$ is large (i.e., if the agglomeration force is small), and unstable otherwise.
Gain functions. Beyond the Beckmann model, the utility gain $\omega$ in Lemma˜1 is well defined for any differentiable $\bm{v}$. In many spatial models, $\omega$ is a simple function of $\Theta$,
$\omega=\Omega(\Theta),$ (6)
as in the Beckmann model. For each model, we call this $\Omega$ the gain function of the model. The positive terms of $\Omega$ represent the model’s agglomeration forces, whereas the negative terms represent the dispersion forces. How each term of $\Omega$ responds to changes in $\Theta$ then describes how transport conditions alter the strengths of these forces, and therefore their relative importance at each transport cost level. [^6] Generalization is possible for the cases where the proximity gain is multi-dimensional such as $\bm{\Theta}\in[0,1]^{M}$ where $M$ is the number of different interregional interactions. For clarity, we restrict our attention to the case $M=1$ throughout this study.